OpenStudy (anonymous):

write an explicit formula for the sequence 4,1.5,-1,-3.5

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

How did I know that was going to happen? :-)

OpenStudy (whpalmer4):

Is this an arithmetic or geometric sequence?

OpenStudy (anonymous):

lol what is the difference

OpenStudy (whpalmer4):

Arithmetic sequence is one where each term differs from the previous term by a fixed amount called the common difference. Geometric sequence is one where each term is the previous term * a fixed amount, called the common ratio. Arithmetic sequence example: 1, 2, 3, 4, 5, 6 ... each one is the previous one + 1 Geometric sequence example: 1, 2, 4, 8, 16, 32 ... each one is the previous one * 2

OpenStudy (anonymous):

oh ok let me see

OpenStudy (anonymous):

im not sure im stuck

OpenStudy (whpalmer4):

Subtract the second term from the first term. Subtract the third term from the second term. Do you get the same answer each time?

OpenStudy (whpalmer4):

4-1.5 = 2.5 1.5 - (-1) = 2.5 -1 - (-3.5) = 2.5 that's an arithmetic sequence. The comparable test for a geometric sequence would be to divide instead of subtracting. For my 1,2,4,8,16 sequence: 2/1 = 2 4/2 = 2 8/4 = 2 etc. that's a geometric sequence.

OpenStudy (anonymous):

ok so it is arithmetic. what do we do next

OpenStudy (whpalmer4):

Well, what is the common difference? What do we subtract from each term to get the next term?

OpenStudy (whpalmer4):

Hello?

OpenStudy (anonymous):

im sorry i got stuck watching the baby i got sidetracked sorry

OpenStudy (anonymous):

oh and 2.5

OpenStudy (whpalmer4):

so we need to write a formula for this sequence. does your book tell you the general formula for an arithmetic sequence?

OpenStudy (whpalmer4):

Usually it is \[a_n = a_1 + (n-1)d\]where \(n\) is the number of the term, and \(d\) is the common difference. We need to use \(d=-2.5\) because each term gets smaller, not larger. What will we use for the value of \(a_1\)?

OpenStudy (whpalmer4):

\(a_1\) is just the first term in the sequence.

OpenStudy (whpalmer4):

The sequence is 4, 1.5, -1, -3.5, ... \[a_1= 4\]\[d = -2.5\]Plug those values into the formula and you have your equation for the sequence.

OpenStudy (whpalmer4):

Test it out by finding the value of \(a_5\) for me.

OpenStudy (whpalmer4):

and \(a_{100}\) also! :-)

OpenStudy (anonymous):

um i dont know how what do i put in for n

OpenStudy (anonymous):

@whpalmer4 im back i will not leave you this time promise. =D

OpenStudy (whpalmer4):

As I said earlier, \(n\) is the number of the term. You're writing a formula, so you won't know the values of everything necessarily. When you are computing \(a_{100}\) you'll use \(n = 100\).

OpenStudy (anonymous):

thats what i thought but when i put a1=4, d=-2.5, and n=5 then answer came out 5.5???

OpenStudy (anonymous):

is it right???

OpenStudy (whpalmer4):

Show me your work...

OpenStudy (anonymous):

\[4+(5-1)-2.5=5.5\] thats what i get when i uses my calculator

OpenStudy (whpalmer4):

\[a_n=a_1+(n−1)d\] \[a_n = 4 + (5-1)(-2.5) =\]

OpenStudy (whpalmer4):

You added -2.5 instead of multiplying by it, I bet! That's what you wrote here, at least...

OpenStudy (anonymous):

ok i see what i did it should be \[4+(5-1)(-2.5)=-6\]

OpenStudy (whpalmer4):

Right. So what does \(a_{100}=\)

OpenStudy (anonymous):

=-243.5

OpenStudy (anonymous):

@whpalmer4 thank you for your help!!

OpenStudy (whpalmer4):

Sorry, fat-fingered the calculation myself, -243.5 is correct :-)

OpenStudy (anonymous):

oh lol i was like what o.0

OpenStudy (whpalmer4):

If you need a handy table of the first 100 terms, here you go :-) \[\begin{array}{cc} 1. & 4. \\ 2. & 1.5 \\ 3. & -1. \\ 4. & -3.5 \\ 5. & -6. \\ 6. & -8.5 \\ 7. & -11. \\ 8. & -13.5 \\ 9. & -16. \\ 10. & -18.5 \\ 11. & -21. \\ 12. & -23.5 \\ 13. & -26. \\ 14. & -28.5 \\ 15. & -31. \\ 16. & -33.5 \\ 17. & -36. \\ 18. & -38.5 \\ 19. & -41. \\ 20. & -43.5 \\ 21. & -46. \\ 22. & -48.5 \\ 23. & -51. \\ 24. & -53.5 \\ 25. & -56. \\ 26. & -58.5 \\ 27. & -61. \\ 28. & -63.5 \\ 29. & -66. \\ 30. & -68.5 \\ 31. & -71. \\ 32. & -73.5 \\ 33. & -76. \\ 34. & -78.5 \\ 35. & -81. \\ 36. & -83.5 \\ 37. & -86. \\ 38. & -88.5 \\ 39. & -91. \\ 40. & -93.5 \\ 41. & -96. \\ 42. & -98.5 \\ 43. & -101. \\ 44. & -103.5 \\ 45. & -106. \\ 46. & -108.5 \\ 47. & -111. \\ 48. & -113.5 \\ 49. & -116. \\ 50. & -118.5 \\ 51. & -121. \\ 52. & -123.5 \\ 53. & -126. \\ 54. & -128.5 \\ 55. & -131. \\ 56. & -133.5 \\ 57. & -136. \\ 58. & -138.5 \\ 59. & -141. \\ 60. & -143.5 \\ 61. & -146. \\ 62. & -148.5 \\ 63. & -151. \\ 64. & -153.5 \\ 65. & -156. \\ 66. & -158.5 \\ 67. & -161. \\ 68. & -163.5 \\ 69. & -166. \\ 70. & -168.5 \\ 71. & -171. \\ 72. & -173.5 \\ 73. & -176. \\ 74. & -178.5 \\ 75. & -181. \\ 76. & -183.5 \\ 77. & -186. \\ 78. & -188.5 \\ 79. & -191. \\ 80. & -193.5 \\ 81. & -196. \\ 82. & -198.5 \\ 83. & -201. \\ 84. & -203.5 \\ 85. & -206. \\ 86. & -208.5 \\ 87. & -211. \\ 88. & -213.5 \\ 89. & -216. \\ 90. & -218.5 \\ 91. & -221. \\ 92. & -223.5 \\ 93. & -226. \\ 94. & -228.5 \\ 95. & -231. \\ 96. & -233.5 \\ 97. & -236. \\ 98. & -238.5 \\ 99. & -241. \\ 100. & -243.5 \\ \end{array}\]

OpenStudy (whpalmer4):

If you need any other values, well, you know the formula :-)

OpenStudy (anonymous):

thanks again